What is the volume of a cube if the are is 169 sq in? (in the book, the area is only given)
Answer by nerdybill(7384)

What is the volume of a cube if the are is 169 sq in? (in the book, the area is only given)
.
We know that the length of all edges of a cube is the same AND that it has 6 sides (faces).
Because it has 6 sides, the area of EACH side (face) of the cube MUST be:
169/6 = 28.1667 sq inch
.
Let x = length of one side of cube
then
x^2 = 28.1667
x = sqrt(28.1667)
x = 5.3072 inches
.
Volume is:
x^3 = 5.3072^3 = 149.487 cubic inches